OpenStudy (anonymous):

Part 1: Simplify f(x) 3x^2-4x+1/3x-1 Part 2: Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f(one-third) is undefined. Be sure to show your work.

OpenStudy (anonymous):

Please don't copy and paste from other sites!! I'm slow and can you show me how you did it and the steps you took! thankyou

OpenStudy (kc_kennylau):

Factorize the numerator first

OpenStudy (kc_kennylau):

Use whatever quadratic factorization method and you'll get 3x^2-4x+1=(3x-1)(x-1)

OpenStudy (kc_kennylau):

And please include parentheses as 3x^2-4x+1/3x-1 is different from (3x^2-4x+1)/(3x-1)

OpenStudy (kc_kennylau):

\[\frac{3x^2-4x+1}{3x-1}=\frac{(3x-1)(x-1)}{3x-1}=?\]

OpenStudy (kc_kennylau):

You should be able to calculate the question mark